OpenStudy (anonymous):

If 16, x, 9 are the first 3 terms of a geometric sequence, find the exact value of x can someone tell me how to find it please :) thanks

OpenStudy (anonymous):

\[{16 \over x}={x \over 9}\]

OpenStudy (anonymous):

\(x=12\)

OpenStudy (anonymous):

\(x=-12\) is also a solution.

OpenStudy (anonymous):

16=4^2 9=3^2 x should be 3.5^2 = 12.25

OpenStudy (anonymous):

A geometric sequence is a sequence in which a term \(a_i\) can be found by multiplying the previous term \(a_{i-1}\) by some constant \(d\). In our case d could be \(-3/4\) or \(3/4\).

OpenStudy (anonymous):

??? i got 3/4 too but thats wrong the right answer is 12 but i dont understand how they got it why did you put it in a fraction???

OpenStudy (anonymous):

Yeah x is either \(12\) or \(-12\). You get it by solving the equation I put in my first comment.

OpenStudy (anonymous):

In a geometric sequence, the quotient of any two consecutive in the sequence must give the same value. So, in our case, \(16/x =x/9\). Solve this equation for x and you will get the answer.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

You're welcome :)